# Conditioning of Banach Space Valued Gaussian Random Variables: An Approximation Approach Based on Martingales

### v1

@@ 5 Proofs of the Results of Section 3
Proof of Theorem 5.10: Let us write ${\cal A}_{n}:=\sigma(A_{n})\subset\sigma(F^{\prime})$ and ${\cal A}_{\infty}:=\sigma(F^{\prime})$. Moreover, we define $A_{\infty}:=\operatorname{id}_{F}$ and $Y_{\infty}:=Y$, which gives $Y_{\infty}=A_{\infty}\circ Y$. Since $(A_{n})$ is a filtering sequence we then have
@@
of $\mathrm{Q}^{(n)}$. Theorem D.2 applied to $\Phi:=\operatorname{id}_{E}$ and $\Psi:=A_{n}$ then gives an $N^{(1)}_{n}\in{\cal A}_{n}$ with $\mathrm{Q}^{(n)}_{F}(N^{(1)}_{n}))=0$ such that for all $y\in F\setminus N^{(1)}_{n}$ and all $B\in\mathfrak{B}(E)$ we have
@@
Let us define
@@
$g=g\circ\pi_{E}(\,\cdot\,,y)\in{{\cal L}_{1}(\mathrm{Q}^{(n)}(\,\cdot\,|y),G)}\,,\qquad\qquad y\in F\setminus N^{(2)}_{n}$
@@
for all $C\in{\cal C}_{n}$. In other words, $Z_{n}\circ\pi_{F}$ is indeed a version of $\mathbb{E}_{\mathrm{P}_{(X,Y)}}(g\circ\pi_{E}|{\cal C}_{n})$. Moreover, these versions are independent of $x\in E$, and for all $y\in F\setminus N_{\mathrm{mart}}$ the martingale convergence (50) thus shows
$Z_{n}(y)\to Z_{\infty}(y)\,.$
Let us define $N_{g}:=N_{\mathrm{trafo}}\cup N_{\mathrm{mart}}\cup N^{(2)}_{\infty}\cup\bigcup_{n\geq 1}N^{(2)}_{n}$. Clearly, this definition gives $N_{g}\in\sigma(F^{\prime})$ with $\mathrm{P}_{Y}(N_{g})=0$. For $y\in F\setminus N_{g}$ our considerations further give
@@
Proof of Theorem 3.4: In the following, we sequentially prove all parts of Theorem 3.4. To this end, we construct for each part x) an $N^{x)}\in\sigma(F^{\prime})$ with $\mathrm{P}_{Y}(N^{x)})=0$ such that for all $y\in F\setminus N^{x)}$ the claims of the parts i) to x) hold true. The desired $\mathrm{P}_{Y}$-zero set $N$ is then given by $N:=N^{viii)}$.
@@
for all $t\in\mathbb{Q}$ and $y\in F\setminus N^{\prime}$ as well as
@@
where the tensor products as well as the space of nuclear operators $\mathfrak{N}(E^{\prime},E)$ are discussed in some detail in front of Lemma B.7 and Lemma B.5, respectively. In addition, we consider the probability space $(E,\mathfrak{B}(E),\mathrm{P}_{X})$ and the function $\tilde{X}:=\operatorname{id}_{E}$. Because $E$ is separable, $\tilde{X}$ is strongly measurable. Since $\mathrm{P}_{X}$ is a Gaussian measure, Fernique’s theorem, see (8), then shows
@@
and hence we find $\tilde{X}\in{{\cal L}_{2}(\mathrm{P}_{X},E)}$. Consequently, Lemma B.7 shows that $g=\tilde{X}\otimes\tilde{X}$ is also strongly measurable with $g\in{{\cal L}_{1}(\mathrm{P}_{X},\mathfrak{N}(E^{\prime},E))}$. Therefore, Theorem 5.10 applied with $G:=\mathfrak{N}(E^{\prime},E)$ shows that there exists an $N_{g}\in\sigma(F^{\prime})$ with $\mathrm{P}_{Y}(N_{g})=0$ such that for all $y\in F\setminus N_{g}$ we have
@@
for all $y\in F\setminus N^{iv)}$. Combining these considerations, we obtain the claimed $C_{X|Y_{n}}\to\operatorname{cov}\mathrm{P}_{X|Y}(\,\cdot\,|y)$ with convergence in $\mathfrak{N}(E^{\prime},E)$. Furthermore, all $C_{X|Y_{n}}$ are independent of $y$, and hence so is their limit $\operatorname{cov}\mathrm{P}_{X|Y}(\,\cdot\,|y)$, that is, the notation $C_{X|Y}$ is justified.
@@
Now, since all $\langle x^{\prime},Z\rangle_{E^{\prime},E}$ are Gaussian, $Z$ is a Gaussian random variable. Moreover, we have $\mathbb{E}\langle x^{\prime},Z_{n}\rangle_{E^{\prime},E}=\langle x^{\prime},\mathbb{E}Z_{n}\rangle_{E^{\prime},E}=\langle x^{\prime},\mathbb{E}X\rangle_{E^{\prime},E}$, and hence (62) gives
$\langle x^{\prime},\mathbb{E}X\rangle_{E^{\prime},E}=\mathbb{E}\langle x^{\prime},Z\rangle_{E^{\prime},E}=\langle x^{\prime},\mathbb{E}Z\rangle_{E^{\prime},E}\,.$
@@
$\langle x^{\prime},Z_{n}-\mathbb{E}Z_{n}\rangle_{E^{\prime},E}\in{\cal G}(Y_{n})\subset{\cal G}(Y)\,.$
Moreover, $Z_{n}\to Z$ in ${{\cal L}_{2}(\mathrm{P},E)}$ implies $\langle x^{\prime},Z_{n}\rangle_{E^{\prime},E}\to\langle x^{\prime},Z\rangle_{E^{\prime},E}$ in ${L_{2}(\mathrm{P})}$. By $\mathbb{E}Z_{n}=\mathbb{E}X=\mathbb{E}Z$ we conclude $\langle x^{\prime},Z_{n}-\mathbb{E}Z_{n}\rangle_{E^{\prime},E}\to\langle x^{\prime},Z-\mathbb{E}Z\rangle_{E^{\prime},E}$ in ${L_{2}(\mathrm{P})}$. Since ${\cal G}(Y)$ is a closed subspace of ${L_{2}(\mathrm{P})}$ by definition, we thus find $\langle x^{\prime},Z-\mathbb{E}Z\rangle_{E^{\prime},E}\in{\cal G}(Y)$. In other words, we have shown ${\cal G}(Z)\subset{\cal G}(Y)$.
To verify the formula for the orthogonal projection, we recall that a bounded linear operator on a Hilbert space is an orthogonal projection if and only if it is idempotent and self-adjoint, see e.g. [11, Proposition II.3.3]. Moreover, $\mathbb{E}(\,\cdot\,|Y):{L_{2}(\mathrm{P})}\to{L_{2}(\mathrm{P})}$ is known to be an orthogonal projection, where we note that this result is rarely stated in the formulation above, so we quickly prove it here using well-known properties of conditional expectations, see e.g. [27, Theorem 8.14]. First, for $f\in{L_{2}(\mathrm{P})}$, we have
@@
$\langle f,\mathbb{E}(g|Y)\rangle_{{L_{2}(\mathrm{P})}}=\mathbb{E}\bigl{(}f\cdot{\mathbb{E}(g|Y)}\bigr{)}=\mathbb{E}\bigl{(}\mathbb{E}(f\cdot{\mathbb{E}(g|Y)}|Y)\bigr{)}$ $=\mathbb{E}\bigl{(}{\mathbb{E}(f|Y)}\cdot{\mathbb{E}(g|Y)}\bigr{)}$
@@
By symmetry we then find
@@
that is, $\mathbb{E}(\,\cdot\,|Y)$ is also self-adjoint. Our next goal is to show
@@
where the first identity holds up to the $\mathrm{P}$-zero set $Y^{-1}(N^{vii)})\in{\cal A}$ and in the second identity we simply considered the version $Z$ of $\mathbb{E}(X|Y)$. Moreover, the third and forth identities are meant in the sense that both $\langle x^{\prime},Z\rangle_{E^{\prime},E}=\langle x^{\prime},\mathbb{E}(X|Y)\rangle_{E^{\prime},E}$ and $\langle x^{\prime},X\rangle_{E^{\prime},E}$ are versions of $\mathbb{E}(\langle x^{\prime},X\rangle_{E^{\prime},E}|Y)$. Consequently, the set
$N_{0}:=\bigl{\{}\langle x^{\prime},Z\rangle_{E^{\prime},E}\neq\langle x^{\prime},X\rangle_{E^{\prime},E}\bigr{\}}$
is a $\mathrm{P}$-zero set with $N_{0}\in\sigma(Y)$. Therefore, there exists an $N^{\prime}\in\sigma(F^{\prime})$ with $Y^{-1}(N^{\prime})=N_{0}$, and in particular we have $\mathrm{P}_{Y}(N^{\prime})=0$. Setting $N^{viii)}:=N^{vii)}\cup N^{\prime}$ then gives the first assertion.
@@
By (B.9) we then find $C_{X|Y}x^{\prime}=0$ as discussed there. ∎
5.6 Proofs of Theorems 3.5 and 3.6
Proof of Theorem 3.5: Throughout the proof we use the shorthand $C_{X|Y}:=\operatorname{cov}(X)-\operatorname{cov}(\mu_{X|Y}\circ Y)$. It suffices to verify the three conditions of Definition 2.5.
@@
Now, the integrand $f(t,y):=\boldsymbol{1}_{A}(t)\exp\bigl{(}-\frac{(t-m(y))^{2}}{2\sigma^{2}}\bigr{)}$ is a non-negative $(\mathfrak{B}\otimes\sigma(F^{\prime}),\mathfrak{B})$-measurable function, and therefore Fubini-Tonelli’s theorem, see e.g. [27, Theorem 14.16], yields the measurability of $y\mapsto{\cal N}(m(y),\sigma^{2})(A)$. In both cases we have thus found $B\in{\cal B}$, and hence we have established ${\cal E}\subset{\cal B}$.
@@
$\lim_{n\to\infty}\bigl{(}\operatorname{cov}(X,Y_{n})(\operatorname{cov}Y_{n})^{\dagger}\operatorname{cov}(Y_{n},X)\bigr{)}(x^{\prime})=\lim_{n\to\infty}\bigl{(}\operatorname{cov}(X,Y_{n})(\operatorname{cov}Y_{n})^{\dagger}\bigr{)}(A_{n}y)$
@@
In particular, $\mu_{X|Y,\mathfrak{A}}$ is the pointwise limit of a sequence of strongly measurable functions $F\to E$, and hence also strongly measurable. Moreover, our construction ensures $\{\mu_{X|Y}\neq\mu_{X|Y,\mathfrak{A}}\}=F_{\mathrm{conv},\mathfrak{A}}\setminus(F\setminus N)$ since we already know $F\setminus N\subset F_{\mathrm{conv},\mathfrak{A}}$. This shows
$\mathrm{P}\bigl{(}\{\mu_{X|Y}\circ Y\neq\mu_{X|Y,\mathfrak{A}}\circ Y\}\bigr{)}=\mathrm{P}_{Y}\bigl{(}\{\mu_{X|Y}\neq\mu_{X|Y,\mathfrak{A}}\}\bigr{)}=\mathrm{P}_{Y}(F_{\mathrm{conv},\mathfrak{A}})-\mathrm{P}_{Y}(F\setminus N)=0\,,$
where we again used $F\setminus N\subset F_{\mathrm{conv},\mathfrak{A}}$. Since $\mu_{X|Y}\circ Y$ is a version of $\mathbb{E}(X|Y)$ by part iv) of Theorem 3.4, we conclude that $\mu_{X|Y,\mathfrak{A}}\circ Y$ is also a version of $\mathbb{E}(X|Y)$.
@@
6 Proofs for the Examples in Section 4
Proof of Proposition 4.1: We will apply Proposition 3.3. To verify (14) we pick a $y^{\prime}\in F^{\prime}$. Without loss of generality we may assume $y^{\prime}\neq 0$. We define $y_{\infty}^{\prime}:=\|y^{\prime}\|^{-1}y^{\prime}$. Since $y_{\infty}^{\prime}\in B_{F^{\prime}}$ and the relative $w^{*}$-topology of $B_{E^{\prime}}$ is metrizable, see Theorem 2.1, there then exists a sequence $(y_{i_{k}}^{\prime})\subset\mathfrak{D}_{F^{\prime}}$ with $y_{i_{k}}^{\prime}\to y_{\infty}^{\prime}$ in the $w^{*}$-topology. This shows

### v3

@@ 6 Proofs of the Results of Section 3
Proof of Theorem 6.7: We begin with some preparations needed for both assertions. To this end, we write ${\cal A}_{n}:=\sigma(A_{n})\subset\sigma(F^{\prime})$ and ${\cal A}_{\infty}:=\sigma(F^{\prime})$. Moreover, we define $A_{\infty}:=\operatorname{id}_{F}$ and $Y_{\infty}:=Y$, which gives $Y_{\infty}=A_{\infty}\circ Y$. Since $(A_{n})$ is a filtering sequence we then have
@@
of $\mathrm{Q}^{(n)}$. Theorem D.2 applied to $\Phi:=\operatorname{id}_{E}:E\to E$ and $\Psi:=A_{n}:F\to\mathbb{R}^{n}$ and the probability measures $\mathrm{P}^{(0)}:=\mathrm{Q}^{(n)}$ and $\mathrm{P}^{(1)}:=\mathrm{P}_{(X,Y_{n})}$ then gives an $N^{(1)}_{n}\in{\cal A}_{n}$ with $\mathrm{Q}^{(n)}_{F}(N^{(1)}_{n}))=0$ such that for all $y\in F\setminus N^{(1)}_{n}$ and all $B\in\mathfrak{B}(E)$ we have
@@
Here we note that in the case $n=\infty$ our construction gives $\mathrm{Q}^{(\infty)}=\mathrm{P}_{(X,Y)}$ as noted above and therefore (49) holds in the case $n=\infty$ for all $y\in F$. Let us now define
@@
${}\displaystyle g=g\circ\pi_{E}(\,\cdot\,,y)\in{{\cal L}_{1}(\mathrm{Q}^{(n)}(\,\cdot\,|y),G)}\,,\qquad\qquad y\in F\setminus N^{(2)}_{n},\,n\in\mathbb{N}\cup\{\infty\}$ (53)
@@
for all $C\in{\cal C}_{n}$. In other words, $Z_{n}\circ\pi_{F}$ is indeed a version of $\mathbb{E}_{\mathrm{P}_{(X,Y)}}(g\circ\pi_{E}|{\cal C}_{n})$. For all $x\in E$ and $y\in F\setminus N_{\mathrm{mart}}$ the martingale convergence (52) thus shows
$Z_{n}(y)=Z_{n}\circ\pi_{F}(x,y)\to Z_{\infty}\circ\pi_{F}(x,y)=Z_{\infty}(y)\,.$
Let us define $N_{g}:=N_{\mathrm{trafo}}\cup N_{\mathrm{mart}}\cup N^{(2)}_{\infty}\cup\bigcup_{n\geq 1}N^{(2)}_{n}$. Clearly, this definition gives $N_{g}\in\sigma(F^{\prime})$ with $\mathrm{P}_{Y}(N_{g})=0$, see the calculation (50). For $y\in F\setminus N_{g}$ our considerations (53) and (49) further give
@@
Proof of Theorem 3.3: In the following, we sequentially prove all parts of Theorem 3.3. To this end, we construct for each part k) an $N^{k)}\in\sigma(F^{\prime})$ with $\mathrm{P}_{Y}(N^{k)})=0$ such that for all $y\in F\setminus N^{k)}$ the claims of the parts i) to k) hold true. The desired $\mathrm{P}_{Y}$-zero set $N$ is then given by $N:=N^{viii)}$.
@@
for all $t\in\mathbb{Q}$ and $y\in F\setminus N^{\prime}$, and since $\mathrm{P}_{X|Y_{n}}(\,\cdot\,|A_{n}(y))={\cal N}(\mu_{n,y}^{(1)},\sigma_{n,y}^{2})$ for such $y$ we also find
@@
where the tensor products as well as the space of nuclear operators $\mathfrak{N}(E^{\prime},E)$ are discussed in some detail in front of Lemma B.7 and Lemma B.5, respectively. In addition, we consider the probability space $(E,\mathfrak{B}(E),\mathrm{P}_{X})$ and the function $\tilde{X}:=\operatorname{id}_{E}$. Because $E$ is separable, $\tilde{X}$ is strongly measurable. In addition, since $\mathrm{P}_{X}$ is a Gaussian measure, Fernique’s theorem, see (8), shows
@@
and hence we find $\tilde{X}\in{{\cal L}_{2}(\mathrm{P}_{X},E)}$. Consequently, Lemma B.7 shows that the $\mathfrak{N}(E,E)$-valued function $g=\tilde{X}\otimes\tilde{X}$ is also strongly measurable with $g\in{{\cal L}_{1}(\mathrm{P}_{X},\mathfrak{N}(E^{\prime},E))}$. Therefore, Theorem 6.7 applied with $G:=\mathfrak{N}(E^{\prime},E)$ shows that there exists an $N_{g}\in\sigma(F^{\prime})$ with $\mathrm{P}_{Y}(N_{g})=0$ such that for all $y\in F\setminus N_{g}$ we have
@@
for all $y\in F\setminus N^{iv)}$. Combining these considerations, we obtain the claimed $C_{X|Y_{n}}\to\operatorname{cov}\mathrm{P}_{X|Y}(\,\cdot\,|y)$ with convergence in $\mathfrak{N}(E^{\prime},E)$. Furthermore, all $C_{X|Y_{n}}$ are independent of $y$, and hence so is their limit $\operatorname{cov}\mathrm{P}_{X|Y}(\,\cdot\,|y)$, that is, the notation $C_{X|Y}$ is indeed justified.
@@
Now, since all $\langle x^{\prime},Z\rangle_{E^{\prime},E}$ are Gaussian, $Z$ is a Gaussian random variable. Moreover, we have $\mathbb{E}\langle x^{\prime},Z_{n}\rangle_{E^{\prime},E}=\langle x^{\prime},\mathbb{E}Z_{n}\rangle_{E^{\prime},E}=\langle x^{\prime},\mathbb{E}X\rangle_{E^{\prime},E}$, and hence we obtain
$\langle x^{\prime},\mathbb{E}X\rangle_{E^{\prime},E}=\lim_{n\to\infty}\langle x^{\prime},\mathbb{E}Z_{n}\rangle_{E^{\prime},E}=\lim_{n\to\infty}\mathbb{E}\langle x^{\prime},Z_{n}\rangle_{E^{\prime},E}=\mathbb{E}\langle x^{\prime},Z\rangle_{E^{\prime},E}=\langle x^{\prime},\mathbb{E}Z\rangle_{E^{\prime},E}\,.$
@@
$\langle x^{\prime},Z_{n}-\mathbb{E}Z_{n}\rangle_{E^{\prime},E}\in{\cal G}(Z_{n})\subset{\cal G}(Y_{n})\subset{\cal G}(Y)\,.$
Moreover, $Z_{n}\to Z$ in ${{\cal L}_{2}(\mathrm{P},E)}$ implies $\langle x^{\prime},Z_{n}\rangle_{E^{\prime},E}\to\langle x^{\prime},Z\rangle_{E^{\prime},E}$ in ${L_{2}(\mathrm{P})}$. By $\mathbb{E}Z_{n}=\mathbb{E}X=\mathbb{E}Z$ we conclude $\langle x^{\prime},Z_{n}-\mathbb{E}Z_{n}\rangle_{E^{\prime},E}\to\langle x^{\prime},Z-\mathbb{E}Z\rangle_{E^{\prime},E}$ in ${L_{2}(\mathrm{P})}$. Since ${\cal G}(Y)$ is a closed subspace of ${L_{2}(\mathrm{P})}$ by definition, we thus find $\langle x^{\prime},Z-\mathbb{E}Z\rangle_{E^{\prime},E}\in{\cal G}(Y)$. Taking the ${L_{2}(\mathrm{P})}$-closure then yields ${\cal G}(Z)\subset{\cal G}(Y)$.
To verify the formula for the orthogonal projection, we recall that a bounded linear operator on a Hilbert space is an orthogonal projection if and only if it is idempotent and self-adjoint, see e.g. [13, Proposition II.3.3]. Moreover, $\mathbb{E}(\,\cdot\,|Y):{L_{2}(\mathrm{P})}\to{L_{2}(\mathrm{P})}$ is known to be an orthogonal projection, where we note that this result is rarely stated in the formulation above, so we quickly prove it here using well-known properties of conditional expectations, see e.g. [29, Theorem 8.14]. First, for $f\in{L_{2}(\mathrm{P})}$, we have
@@
$\langle f,\mathbb{E}(g|Y)\rangle_{{L_{2}(\mathrm{P})}}=\mathbb{E}\bigl{(}f\cdot{\mathbb{E}(g|Y)}\bigr{)}=\mathbb{E}\bigl{(}\mathbb{E}\bigl{(}f\cdot{\mathbb{E}(g|Y)}|Y\bigr{)}\bigr{)}$ $=\mathbb{E}\bigl{(}{\mathbb{E}(f|Y)}\cdot{\mathbb{E}(g|Y)}\bigr{)}$
@@
Analogously, we obtain $\langle g,\mathbb{E}(f|Y)\rangle_{{L_{2}(\mathrm{P})}}=\bigl{\langle}\mathbb{E}(f|Y),\mathbb{E}(g|Y)\bigr{\rangle}_{{L_{2}(\mathrm{P})}}$, and thus we also find
@@
that is, $\mathbb{E}(\,\cdot\,|Y)$ is also self-adjoint. In summary, $\mathbb{E}(\,\cdot\,|Y):{L_{2}(\mathrm{P})}\to{L_{2}(\mathrm{P})}$ is an orthogonal projection. Our next goal is to show
@@
where the first identity holds up to the $\mathrm{P}$-zero set $Y^{-1}(N^{vii)})\in{\cal A}$ and in the second identity we simply considered the version $Z$ of $\mathbb{E}(X|Y)$. Moreover, the third and forth identities are meant in the sense that both $\langle x^{\prime},\mathbb{E}(X|Y)\rangle_{E^{\prime},E}$ and $\langle x^{\prime},X\rangle_{E^{\prime},E}$ are versions of $\mathbb{E}(\langle x^{\prime},X\rangle_{E^{\prime},E}|Y)$. Consequently, the set
$N_{x^{\prime}}:=\bigl{\{}\langle x^{\prime},\mu_{X|Y}\circ Y\rangle_{E^{\prime},E}\neq\langle x^{\prime},X\rangle_{E^{\prime},E}\bigr{\}}$
is a $\mathrm{P}$-zero set with $N_{x^{\prime}}\in\sigma(Y)$. Let us now consider the set
[…]
as well as a metric $\operatorname{d}_{w^{*}}$ describing the relative $w^{*}$-topology of $B_{E^{\prime}}$. Since subsets of separable metric spaces are separable, and $(B_{E^{\prime}},\operatorname{d}_{w^{*}})$ is separable by Theorem 2.1, we conclude that there exists a countable ${\cal F}_{0}\subset{\cal F}$ that is $\operatorname{d}_{w^{*}}$-dense in ${\cal F}$. We define
$N_{{\cal F}_{0}}:=\bigcup_{x^{\prime}\in{\cal F}_{0}}N_{x^{\prime}}\,.$
By construction, we have $N_{{\cal F}_{0}}\in\sigma(Y)$ with $\mathrm{P}(N_{{\cal F}_{0}})=0$. Moreover, $N_{{\cal F}_{0}}\in\sigma(Y)$ gives an $N_{{\cal F}_{0}}^{\prime}\in\sigma(F^{\prime})$ with $Y^{-1}(N_{{\cal F}_{0}}^{\prime})=N_{{\cal F}_{0}}$ and $\mathrm{P}_{Y}(N_{{\cal F}_{0}}^{\prime})=0$. We define $N^{viii)}:=N^{vii)}\cup N_{{\cal F}_{0}}^{\prime}$. By construction, we then have
$\bigl{\langle}x^{\prime},\mu_{X|Y}(Y(\omega))\bigr{\rangle}_{E^{\prime},E}=\langle x^{\prime},X(\omega)\rangle_{E^{\prime},E}\,,\qquad\qquad x^{\prime}\in{\cal F}_{0},\omega\not\in Y^{-1}(N^{viii)})\,.$
Since ${\cal F}_{0}$ is $\operatorname{d}_{w^{*}}$-dense in ${\cal F}$, we then find the first assertion for all $x^{\prime}\in{\cal F}$, and a simple rescaling argument then gives the desired identity in the general case.
@@
By (B.9) this implies $C_{X|Y}x^{\prime}=0$ as discussed there. ∎
6.6 Proofs of Theorems 3.4 and 3.5
Proof of Theorem 3.4: Throughout the proof we use the shorthand $C_{X|Y}:=\operatorname{cov}(X)-\operatorname{cov}(\mu\circ Y)$. It suffices to verify the three conditions of Definition 2.5.
@@
Now, the integrand $f(t,y):=\boldsymbol{1}_{A}(t)\exp\bigl{(}-\frac{(t-m(y))^{2}}{2\sigma^{2}}\bigr{)}$ is a non-negative $(\mathfrak{B}\otimes\sigma(F^{\prime}),\mathfrak{B})$-measurable function, and therefore Fubini-Tonelli’s theorem, see e.g. [29, Theorem 14.16], yields the measurability of $y\mapsto{\cal N}(m(y),\sigma^{2})(A)$. In both cases we have thus found $B\in{\cal B}$, and hence we have established ${\cal E}\subset{\cal B}$.
@@
$\lim_{n\to\infty}\bigl{(}\operatorname{cov}(X,Y_{n})(\operatorname{cov}Y_{n})^{\dagger}\operatorname{cov}(Y_{n},X)\bigr{)}(x^{\prime})=\lim_{n\to\infty}\bigl{(}\operatorname{cov}(X,Y_{n})(\operatorname{cov}Y_{n})^{\dagger}\bigr{)}(A_{n}y)=\lim_{n\to\infty}B_{n}y$
@@
In particular, $\mu_{X|Y,\mathfrak{A}}$ is the pointwise limit of a sequence of strongly measurable functions $F\to E$, and hence also strongly measurable. Moreover, our construction ensures $F_{\mathrm{conv},\mathfrak{A}}\cap F\setminus N\subset\{\mu_{X|Y}=\mu_{X|Y,\mathfrak{A}}\}$, and since we already know $F\setminus N\subset F_{\mathrm{conv},\mathfrak{A}}$ we conclude that $\{\mu_{X|Y}\neq\mu_{X|Y,\mathfrak{A}}\}\subset N$. This shows
$\mathrm{P}\bigl{(}\{\mu_{X|Y}\circ Y\neq\mu_{X|Y,\mathfrak{A}}\circ Y\}\bigr{)}=\mathrm{P}_{Y}\bigl{(}\{\mu_{X|Y}\neq\mu_{X|Y,\mathfrak{A}}\}\bigr{)}\leq\mathrm{P}_{Y}(N)=0\,.$
Since $\mu_{X|Y}\circ Y$ is a version of $\mathbb{E}(X|Y)$ by part iv) of Theorem 3.3, we conclude that $\mu_{X|Y,\mathfrak{A}}\circ Y$ is also a version of $\mathbb{E}(X|Y)$.
@@
7 Proofs Related to Filtering Sequences in Section 4
This section contains all proofs for the results of Section 4. In addition, some auxiliary results are presented. We begin with the following lemma that considers the behavior of generated $\sigma$-algebras under some natural operations in $F^{\prime}$.
Lemma 7.1.
Let $F$ be a Banach space and $B\subset F^{\prime}$ be non-empty. Then we have
Proof of Lemma 7.1: We first prove the equation $\sigma(B)=\sigma(\operatorname{span}B)$. Here, the inclusion $\subset$” is obvious. To establish the converse, we fix an $y^{\prime}\in\operatorname{span}B$. Then there exist an $n\geq 1$ as well as $\alpha_{1},\dots,\alpha_{n}\in\mathbb{R}$ and $y_{1}^{\prime},\dots,y_{n}^{\prime}\in B$ with $y^{\prime}=\alpha_{1}y_{1}^{\prime}+\dots+\alpha_{n}y_{n}^{\prime}$. Clearly, each $y_{i}^{\prime}:F\to\mathbb{R}$ is $(\sigma(B),\mathfrak{B})$-measurable, and hence $y^{\prime}$ is also $(\sigma(B),\mathfrak{B})$-measurable. This gives $\sigma(y^{\prime})\subset\sigma(B)$, and since $y^{\prime}\in\operatorname{span}B$ was chosen arbitrarily, we find $\sigma(\operatorname{span}B)\subset\sigma(B)$.
The third equation is a direct consequence of the first two equations applied to $B$ and $\operatorname{span}B$. ∎
Lemma 7.2.
  1. i)
  2. ii)
Proof of Lemma 7.2: i). We first note that in view of Lemma 7.1 we only need to show $\sigma(A)=\sigma(\mathfrak{C}_{A})$. To establish this identity, let $a_{1}^{\prime},\dots,a_{n}^{\prime}:F\to\mathbb{R}$ be the coordinate functions of $A$, that is $\mathfrak{C}_{A}=\{a_{1}^{\prime},\dots,a_{n}^{\prime}\}$. Then, [29, Theorem 1.90] shows that for a given $\sigma$-algebra ${\cal A}^{\prime}$ on $F$, the map $A:F\to\mathbb{R}^{n}$ is $({\cal A}^{\prime},\mathfrak{B}^{n})$-measurable if and only if all $a_{1}^{\prime},\dots,a_{n}^{\prime}$ are $({\cal A}^{\prime},\mathfrak{B})$-measurable.
ii). Let $e_{1},\dots,e_{n}$ be the standard ONB of $\mathbb{R}^{n}$. Then $A$ is given by $A=\sum_{i=1}^{n}\langle a_{i}^{\prime},\,\cdot\,\rangle_{F^{\prime},F}e_{i}=\sum_{i=1}^{n}a_{i}^{\prime}\otimes e_{i}$, see (B.14), and therefore (B.16) shows that its adjoint $A^{\prime}$ is given by
where the first identity can be found in e.g. [34, Lemma 3.1.16] or [13, Proposition VI.1.8], and the second identity can be found in e.g. [34, Proposition 2.6.6]. Moreover, since the subspace $\operatorname{ran}A^{\prime}$ of $F^{\prime}$ is finite dimensional, it is $w^{*}$-closed, see e.g. [34, Corollary 2.2.32] or [40, Theorem 1.21]. Combining the latter with (70) yields $(\ker A)^{\perp\perp}=\operatorname{ran}A^{\prime}$. ∎
Proposition 7.3.
$F_{\infty}^{\prime}:=\bigcup_{n\geq 1}\operatorname{span}\mathfrak{C}_{A_{n}}=\bigcup_{n\geq 1}F_{n}^{\prime}\,.$
$=\sigma\biggl{(}\bigcup_{n\geq 1}\sigma(A_{n})\biggr{)}\,,$
Applying Proposition 7.3 then yields the assertion. ∎
Proof of Theorem 4.2: We will apply Proposition 4.1. To verify (19) we pick a $y^{\prime}\in F^{\prime}$. Without loss of generality we may assume $y^{\prime}\neq 0$. We define $y_{\infty}^{\prime}:=\|y^{\prime}\|^{-1}y^{\prime}$. Since $y_{\infty}^{\prime}\in B_{F^{\prime}}$ and the relative $w^{*}$-topology of $B_{E^{\prime}}$ is metrizable, see Theorem 2.1, there then exists a sequence $(y_{i_{k}}^{\prime})\subset\mathfrak{D}_{F^{\prime}}$ with $y_{i_{k}}^{\prime}\to y_{\infty}^{\prime}$ in the $w^{*}$-topology. This shows